# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v3

@@ 5 Optimizely Example
Optimizely is a company that provides software as a service. Their software has helped thousands of other businesses – including StubHub, Starbucks, eBay, Visa, and The New York Times – conduct online controlled experiments (Siroker and Koomen, 2013). Optimizely also advocates for improving their own online experiences; Siroker and Koomen (2013) described an A/B test that Optimizely ran amidst a 2012 redesign of their website. In this section, we illustrate the use of our methods to design an experiment with $K=5$ of the metrics considered by Optimizely during their redesign. Each metric corresponds to a distinct binary outcome for visitor $i=1,\dots,n$:
@@
This example is interesting because these five binary outcomes are not independent. In particular, a visitor cannot experience the final four outcomes unless they engage with the experiment. Moreover, the account creation dialog related to the fourth outcome must be shown for a visitor to successfully create an Optimizely account. These dependent outcomes give rise to a group of hypothesis tests that are not independent as detailed later. While this dependence structure illustrates our method can be applied with complex designs, we emphasize that our method can be broadly applied to design A/B tests. We do not require a complicated dependence structure between the outcomes, and the outcomes can be of varied data types.
We circumvent the need to directly specify the dependence structure between the Optimizely outcomes by jointly modeling all five binary variables using a multinomial model. Figure 1 visualizes this multinomial model, which has 13 probabilities associated with the five outcomes. We first introduce the model for a single treatment group and later introduce subscripts for the group $j$, where group A is the original site and group B is the redesign. In a particular group, the marginal probability of a visitor attaining the $k^{\text{th}}$ binary outcome is represented by $\pi_{k},~{}k=1,\dots,5$. These marginal probabilities can be expressed as sums of various multinomial probabilities $\{\eta_{v}\}_{v=1}^{13}$ from Figure 1: $\pi_{1}=1-\eta_{1}$, $\pi_{2}=\sum_{v=8}^{13}\eta_{v}$, $\pi_{3}=\sum_{v=5}^{7}\eta_{v}+\sum_{v=11}^{13}\eta_{v}$, $\pi_{4}=\sum_{v=1}^{4}\eta_{3v}+\sum_{v=1}^{4}\eta_{3v+1}$, $\pi_{5}=\sum_{v=1}^{4}\eta_{3v+1}$.
@@
For the $k^{\text{th}}$ binary outcome, Siroker and Koomen (2013) considered the relative difference between the probabilities of a visitor attaining that outcome in group B vs. group A. This metric is often called lift in the context of A/B tests. We use lift as the metric of interest for this example:
@@
where the first subscript now denotes the group $j=A,B$. Our hypotheses of interest are $H_{0,k}:\theta_{k}\leq 0$ vs. $H_{1,k}:\theta_{k}>0$. That is, $(\delta_{L,k},\delta_{U,k})=(0,\infty)$ for all metrics $k=1,\dots,5$. This experiment therefore aims to demonstrate that the redesigned website is superior to the original one with respect to the various binary outcomes. The notion of practically important effect sizes could be accounted for using alternative choices for $\delta_{L,k}$. We reiterate that these hypotheses $\{H_{1,k}\}_{k=1}^{5}$ are not independent due to the relationship between $\{\pi_{j,k}\}_{k=1}^{5}$ and $\boldsymbol{\eta}_{j}=\{\eta_{j,v}\}_{v=1}^{13}$.
@@
To implement our method, we must define the probability model $\Psi$ that characterizes data generation. We specify this model by choosing various sets of multinomial models. Each set of models defines separate multinomial probabilities for the two groups. Specifying various sets of multinomial models allows us to accommodate uncertainty in which hypotheses $H_{1,k}$ are false. It also allows us to account for uncertainty in effect sizes for a given set of true hypotheses, but we do not utilize that flexibility in this example. For all sets of multinomial models we consider, the marginal probabilities for group A align with the estimated marginal probabilities for the original site reported by Siroker and Koomen (2013): $\hat{\pi}_{A,1}=0.489,\hat{\pi}_{A,2}=0.230,\hat{\pi}_{A,3}=0.156,\hat{\pi}_{A,4}=0.047,\hat{\pi}_{A,5}=0.032$. Optimizely only provided the marginal probabilities, but we must specify the 13 multinomial probabilities $\{\eta_{A,v}\}_{v=1}^{13}$ for group A. Below, we overview a method to choose these multinomial probabilities and those for group B, and we note that businesses should have historical information about dependence between the outcomes in practice.
@@
We choose each set of multinomial probabilities for both groups using linear programming (Guenin et al., 2014). Linear programming allows us to construct multinomial models that satisfy the following objectives. First, the probabilities $\sum_{v=1}^{13}\eta_{j,v}$ must sum to 1 in both groups. Second, we must attain specified values for the marginal probabilities $\{\pi_{A,k}\}_{k=1}^{5}$ in group A. Third, we need the marginal probabilities $\{\pi_{B,k}\}_{k=1}^{5}$ in group B to achieve targets for lift based on anticipated values for $\{\theta_{k}\}_{k=1}^{5}$. These anticipated values determine which hypotheses in $\{H_{1,k}\}_{k=1}^{5}$ are true along with the effect sizes corresponding to the true hypotheses. Linear programs can also accommodate additional constraints (e.g., enforcing a lower bound for all probabilities in the multinomial model). We elaborate on this method to select a set of multinomial models in Appendix C of the supplement. User-friendly functions to implement this method with linear programming and Algorithm 2 are made available in this GitHub repository: https://github.com/lmhagar/ABTestFDR. These functions can be applied to design A/B tests with an arbitrary number of binary outcomes – regardless of whether they are dependent or independent.
@@
For a general A/B test, there are ($K\choose k$ ways to classify $k$ of the $K$ hypotheses as false. In this Optimizely experiment, there are $\sum_{k=1}^{4}{5\choose k}=30$ unique classifications of $\{H_{1,k}\}_{k=1}^{5}$ such that some but not all of the hypotheses are false. We choose to define $\Psi$ by specifying 30 sets of multinomial models in Appendix C. We index these submodels by introducing a subscript to the notation $\Psi_{\boldsymbol{\Lambda}}$, where $\varnothing\subsetneq\boldsymbol{\Lambda}\subsetneq\{1,2,3,4,5\}$ is the set of hypotheses that are false. For the submodel $\Psi_{\boldsymbol{\Lambda}}$, the multinomial models for each group ensure that $\theta_{k}=0$ for $k\in\boldsymbol{\Lambda}$ and $\theta_{k}=10\%$ otherwise. It follows that $H_{1,k}$ is false in this submodel if and only if $k\in\boldsymbol{\Lambda}$. We define the model $\Psi$ input into Algorithm 2 as an equally weighted mixture of these 30 submodels. While not explored in this example, $\Psi$ could instead be an unequally weighted mixture of submodels.
@@
We now apply the method in Algorithm 2 with balanced sample sizes ($c=1$) along with criteria for the FDR and average power defined by $q=0.05$ and $\beta=0.2$. We choose an initial sample size of $n_{0}=1.2\times 10^{4}$ using information from Optimizely’s 2012 experiment (Siroker and Koomen, 2013). We implemented our method with $10^{3}$ simulation repetitions for each of the 30 submodels, giving rise to a total of $m=3\times 10^{4}$ simulation repetitions. The prior $p(\boldsymbol{\eta})$ is comprised of independent, diffuse Dirichlet$(\mathbf{1}_{13})$ priors for the multinomial probabilities $\boldsymbol{\eta}_{j}$ in each group. These distributions are conjugate priors for $\boldsymbol{\eta}_{A}$ and $\boldsymbol{\eta}_{B}$, which we use to estimate the joint posterior of $\boldsymbol{\theta}$ using simulation.
@@
Algorithm 2 with these inputs returned a recommended $(n_{A},\gamma)$ combination of (13328,(13328,( 13328 , 0.9411)0.9411)0.9411 ). All sample size recommendations in this section pertain to the sample size per group (i.e., $n_{A}=n_{B}$). It took roughly 3 minutes to obtain this recommendation on a parallel computing server with 72 cores. While each posterior takes less than one second to approximate, we must approximate many posteriors to reliably characterize the uncertainty regarding which hypotheses are false. Our method could be expedited by considering multivariate analytical approximations to the posterior of $\boldsymbol{\theta}$. The 95% bootstrap confidence intervals for $n_{A}$ and $\gamma$ obtained with $M=5\times 10^{3}$ bootstrap samples were respectively $[13108,13624]$ and $(0.9398,0.9431)$. To confirm the strong performance of our method, we estimated the joint sampling distribution of posterior probabilities for $n_{A}=13328$ under the model $\Psi$ with $m=9.9\times 10^{4}$ simulation repetitions. Using a decision threshold of $\gamma=0.9411$, the FDR and average power in (1) were estimated as 0.0497 and 0.7992.
@@
To demonstrate the flexibility of our method with different models and the value of considering the dependence structure at the design stage, we implemented Algorithm 2 with a modified version of $\Psi$ where the 5 five binary outcomes are generated independently according to the marginal probabilities from a given submodel $\Psi_{\boldsymbol{\Lambda}}$. The individual metrics in $\boldsymbol{\theta}$ were also analyzed using posteriors based on independent binomial models for each outcome and diffuse BETA(1,1) priors for each marginal probability. When ignoring the dependence structure between the 5 metrics, Algorithm 2 returned a recommended $(n_{A},\gamma)$ combination of $(14427,0.9500)$. For this $(n_{A},\gamma)$ combination with $m=9.9\times 10^{4}$ simulation repetitions, we confirmed the FDR and average power under the binomial data generation process to be 0.0502 and 0.7992. The 95% bootstrap confidence intervals for $n_{A}$ and $\gamma$ in this scenario were respectively $[14138,14742]$ and $(0.9485,0.9516)$. The precision of these bootstrap confidence intervals suggests that the $(n_{A},\gamma)$ recommendation is materially different when accounting for the dependence structure. When accounting for the positive dependence between these binary outcomes, we can use a less strict decision threshold, resulting in a smaller sample size recommendation. If the outcomes exhibit negative dependence, we generally need a larger decision threshold – and sample size – to satisfy criteria for both the FDR and average power.
@@
To illustrate how the number of false hypotheses impacts the sample size recommendation, we implement Algorithm 2 where $\Psi$ is defined using only some of the 30 multinomial submodels. We first consider the $(n_{A},\gamma)$ recommendation where $\Psi$ is an equal mixture of the 5 submodels where only 1 hypothesis is false (i.e., where the cardinality $\lVert\boldsymbol{\Lambda}\rVert$ of the set $\boldsymbol{\Lambda}$ is 1). We separately obtain $(n_{A},\gamma)$ recommendations for the 10 submodels where $\lVert\boldsymbol{\Lambda}\rVert=2$, the 10 submodels where $\lVert\boldsymbol{\Lambda}\rVert=3$, and the 5 submodels where $\lVert\boldsymbol{\Lambda}\rVert=4$. Each recommendation was obtained via Algorithm 2 with $m=3\times 10^{4}$, $q=0.05$, and $\beta=0.2$. 95% bootstrap confidence intervals for each recommendation were computed with $M=5\times 10^{3}$. In Table 1, these results are compared to the $(n_{A},\gamma)$ recommendation where $\Psi$ was defined by combining all 30 multinomial submodels in an equal mixture.
$\Psi$ $n_{A}$ $\gamma$
$\lVert\boldsymbol{\Lambda}\rVert=1$ 4627 [4468, 4761] 0.7772 (0.7718, 0.7819)
$\lVert\boldsymbol{\Lambda}\rVert=2$ 9985 [9693, 10065] 0.9053 (0.9031, 0.9077)
$\lVert\boldsymbol{\Lambda}\rVert=3$ 15035 [14755, 15272] 0.9516 (0.9499, 0.9532)
$\lVert\boldsymbol{\Lambda}\rVert=4$ 19668 [19369, 20024] 0.9737 (0.9730, 0.9745)
Combined 13328 [13108, 13624] 0.9411 (0.9398, 0.9431)
@@
Table 1: Recommended sample sizes $n_{A}$ and decision thresholds $\gamma$ under probability models $\Psi$ that do (Combined) and do not ($\lVert\boldsymbol{\Lambda}\rVert=\{1,2,3,4\}$) account for uncertainty in the number of false hypotheses.
@@
Lastly, we compare our sample size recommendations to those prompted by a Bonferroni correction used to control the FWER of A/B tests in frequentist settings (Kohavi et al., 2020). To bound the FWER for 5 hypotheses by 0.05, the Bonferroni correction recommends using a significance level of 0.01 for each test. This significance level is similar to a decision threshold of $\gamma=0.99$ in our context. We first separately consider the sample sizes per group required to detect 10% lift with 80% power for each of the hypotheses in $\{H_{1},k\}_{k=1}^{5}$. Given the marginal probabilities $\{\pi_{A,k}\}_{k=1}^{5}$ in group A, these sample sizes $n_{A}$ for $\{H_{1,k}\}_{k=1}^{5}$ are respectively 2104, 6961, 11311, 42667, and 63702. These sample size recommendations differ because the magnitude of the difference between the marginal probabilities $\pi_{B,k}$ and $\pi_{A,k}$ that define 10% lift depends on the metric $k$. A sample size of $n_{A}=27956$ is recommended to ensure the average power across these 5 hypothesis tests is at least 80% while bounding the FWER by 0.05. We emphasize that control of the FDR and FWER are not interchangeable, but by comparing these sample size recommendations to those in Table 1, we nevertheless illustrate that practitioners can reduce the required sample size for their A/B tests if willing to use designs that bound the FDR instead of the FWER.
6 Contour Plots to Explore Design Criteria
Algorithm 2 returns the $(n,\gamma)$ combination with the smallest sample size $n$ that satisfies the criteria for both operating characteristics in (1), but practitioners may want to explore designs that are similar to the optimal one. We obtained linear approximations $\hat{L}^{{}_{(n)}}_{r,k}$ for each metric $k$ in Line 12 of Algorithm 2 using estimates of the sampling distributions of posterior probabilities under $\Psi$ at two sample sizes: $n_{0}$ and $n_{1}$. We approximate the joint sampling distribution of posterior probabilities for other sample sizes using the functions $\{\{\hat{L}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$. We use contour plots to synthesize these approximations to the sampling distribution. These plots visualize how changes to $n$ and the decision threshold $\gamma$ impact average power and the FDR.
Figure 2: Left: Contour plots for the FDR and average power from the sample size calculation in Section 5 with the optimal $(n_{A},\gamma)$ combination in gray. Right: Contour plots obtained by estimating sampling distributions throughout the $n_{A}$-space.
The left column of Figure 2 depicts the contour plots with respect to the FDR and average power for the sample size calculation in Section 5 where $\Psi$ was defined using an equally weighted mixture of 30 multinomial models. To assist with interpretation, the green contour corresponding to average power of $0.81-\beta=$ 0.8 and the red contour corresponding to an FDR of $q=0.05$ are depicted on both plots. The criteria for average power and the FDR are respectively satisfied for the regions of the $(n_{A},\gamma)$-space that are below the green contour and above the red contour. The estimated optimal $(n_{A},\gamma)$ combination of $(13328,0.9411)$ is depicted by the gray point. The estimated optimal sample size of $n_{A}=13328$ is the smallest $n_{A}\in\mathbb{Z}^{+}$ that is to the right of the intersection of the red and green contours. The left contour plots simply repurposed the logits of posterior probabilities computed in Algorithm 2. Thus, these plots can be generated in about the same amount of time that it takes to implement Algorithm 2 (roughly 3 minutes on a parallel computing server for this example). These contour plots allow practitioners to visualize the trade-off between the FDR and average power for various sample sizes and decision thresholds without implementing extensive numerical studies.
The contour plots in the right column of Figure 2 are confirmatory and meant to assess the accuracy of those in the left column. They were created by simulating $m=9.9\times 10^{4}$ samples according to $\Psi$ for $n_{A}=\{1\times 10^{4},1.05\times 10^{4},\dots,1.5\times 10^{4}\}$ using the process detailed in Algorithm 1. Even when only considering sample sizes in increments of 500, this process took about 50 minutes using a parallel computing server with 72 cores. The smallest $n_{A}\in\mathbb{Z}^{+}$ to the right of the intersection of the green and red contours in the right plots is $n_{A}=13323$, which is very close to the estimated optimal sample size of $n_{A}=13328$ from Section 5. Moreover, the plots in the left and right columns are similar, which is a consequence of Theorem 1. This similarity builds confidence in our method to construct contour plots by repurposing posterior probabilities from Algorithm 2.

### v4

@@ 5 Optimizely Example
Optimizely is a company that provides software to help other businesses conduct online controlled experiments. Optimizely also advocates for improving their own online experiences; Siroker and Koomen (2013) described an A/B test that Optimizely ran amidst a 2012 redesign of their website. In this section, we illustrate the use of our methods to design an experiment with $K=5$ of the metrics considered by Optimizely during their redesign. The number of metrics we consider ($K=5$) is typical for A/B tests (Ghosh, 2021) and aligns with the rough rule of thumb provided by Kohavi et al. (2020). Each metric corresponds to a distinct binary outcome for visitor $i=1,\dots,n$:
@@
This example is interesting because these five binary outcomes are not independent. In particular, a visitor cannot experience the final four outcomes unless they engage with the experiment. Moreover, the account creation dialog related to the fourth outcome must be shown for a visitor to successfully create an Optimizely account. We circumvent the need to directly specify the dependence structure between the Optimizely outcomes by jointly modeling all five binary variables using a multinomial model. Figure 1 visualizes this multinomial model, which has 13 probabilities associated with the five outcomes. We first introduce the model for a single treatment group and later introduce subscripts for the group $j$, where group A is the original site and group B is the redesign. In a particular group, the marginal probability of a visitor attaining the $k^{\text{th}}$ binary outcome is represented by $\pi_{k},~{}k=1,\dots,5$. These marginal probabilities can be expressed as sums of various multinomial probabilities $\{\eta_{v}\}_{v=1}^{13}$ from Figure 1: $\pi_{1}=1-\eta_{1}$, $\pi_{2}=\sum_{v=8}^{13}\eta_{v}$, $\pi_{3}=\sum_{v=5}^{7}\eta_{v}+\sum_{v=11}^{13}\eta_{v}$, $\pi_{4}=\sum_{v=1}^{4}\eta_{3v}+\sum_{v=1}^{4}\eta_{3v+1}$, $\pi_{5}=\sum_{v=1}^{4}\eta_{3v+1}$.
@@
For the $k^{\text{th}}$ binary outcome, Siroker and Koomen (2013) considered the relative difference (i.e., the lift) between the probabilities of a visitor attaining that outcome in group B vs. group A. We use lift as the metric of interest for this example:
@@
where the first subscript now denotes the group $j=A,B$. Our hypotheses of interest are $H_{0,k}:\theta_{k}\leq 0$ vs. $H_{1,k}:\theta_{k}>0$. That is, $(\delta_{L,k},\delta_{U,k})=(0,\infty)$ for all metrics $k=1,\dots,5$. Thus, this experiment aims to demonstrate that the redesigned website is superior to the original one with respect to the various binary outcomes. We emphasize that the hypotheses $\{H_{1,k}\}_{k=1}^{5}$ are not independent due to the relationship between $\{\pi_{j,k}\}_{k=1}^{5}$ and $\boldsymbol{\eta}_{j}=\{\eta_{j,v}\}_{v=1}^{13}$.
@@
To implement our method, we must define the probability model $\Psi$ that characterizes data generation. We specify this model by choosing various sets of multinomial models. Each set of models defines separate multinomial probabilities for the two groups. Specifying various sets of multinomial models allows us to accommodate uncertainty in which hypotheses $H_{1,k}$ are false. It also allows us to account for uncertainty in effect sizes for a given set of true hypotheses, but we do not utilize that flexibility in this example. For all sets of multinomial models we consider, the marginal probabilities for group A align with the estimated marginal probabilities for the original site reported by Siroker and Koomen (2013): $\hat{\pi}_{A,1}=0.489,\hat{\pi}_{A,2}=0.230,\hat{\pi}_{A,3}=0.156,\hat{\pi}_{A,4}=0.047,\hat{\pi}_{A,5}=0.032$. Optimizely only provided the marginal probabilities, but we must specify the 13 multinomial probabilities $\{\eta_{A,v}\}_{v=1}^{13}$ for group A. Below, we overview a method to choose these multinomial probabilities and those for group B.
@@
We choose each set of multinomial probabilities for both groups using linear programming (Guenin et al., 2014). Linear programming allows us to construct multinomial models that satisfy the following objectives. First, the probabilities $\sum_{v=1}^{13}\eta_{j,v}$ must sum to 1 in both groups. Second, we must attain specified values for the marginal probabilities $\{\pi_{A,k}\}_{k=1}^{5}$ in group A. Third, we need the marginal probabilities $\{\pi_{B,k}\}_{k=1}^{5}$ in group B to achieve targets for lift based on anticipated values for $\{\theta_{k}\}_{k=1}^{5}$. These anticipated values determine which hypotheses in $\{H_{1,k}\}_{k=1}^{5}$ are true along with the effect sizes corresponding to the true hypotheses. We elaborate on this method to select a set of multinomial models in Appendix C.1 of the supplement.
@@
For a general A/B test, there are ($K\choose k$ ways to classify $k$ of the $K$ hypotheses as false. In this Optimizely experiment, there are $\sum_{k=1}^{4}{5\choose k}=30$ unique classifications of $\{H_{1,k}\}_{k=1}^{5}$ such that some but not all of the hypotheses are false. We choose to define $\Psi$ by specifying 30 sets of multinomial models in Appendix C.1. We index these submodels by introducing a subscript to the notation $\Psi_{\boldsymbol{\Lambda}}$, where $\varnothing\subsetneq\boldsymbol{\Lambda}\subsetneq\{1,2,3,4,5\}$ is the set of hypotheses that are false. For the submodel $\Psi_{\boldsymbol{\Lambda}}$, the multinomial models for each group ensure that $\theta_{k}=0$ for $k\in\boldsymbol{\Lambda}$ and $\theta_{k}=10\%$ otherwise. It follows that $H_{1,k}$ is false in this submodel if and only if $k\in\boldsymbol{\Lambda}$. For illustration, we define the model $\Psi$ input into Algorithm 2 as an equally weighted mixture of these 30 submodels.
@@
We now apply the method in Algorithm 2 with balanced sample sizes ($c=1$) along with criteria for the FDR and average power defined by $q=0.05$ and $\beta=0.2$. For now, we find optimal decision thresholds under the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. This constraint gives rise to a simple optimization scheme to select $\gamma$: the optimal decision threshold is the smallest probability such that $\widehat{\text{FDR}}\leq q$. We choose an initial sample size of $n_{0}=1.2\times 10^{4}$ using information from Optimizely’s 2012 experiment (Siroker and Koomen, 2013). We implemented our method with $10^{3}$ simulation repetitions for each of the 30 submodels, giving rise to a total of $m=3\times 10^{4}$ simulation repetitions. The prior $p(\boldsymbol{\eta})$ is comprised of independent, diffuse Dirichlet$(\mathbf{1}_{13})$ priors for the multinomial probabilities $\boldsymbol{\eta}_{j}$ in each group. These distributions are conjugate priors for $\boldsymbol{\eta}_{A}$ and $\boldsymbol{\eta}_{B}$, which we use to estimate the joint posterior of $\boldsymbol{\theta}$ via simulation.
@@
Algorithm 2 with these inputs returned a recommended $(n_{A},\gamma)$ combination of (13328,(13328,( 13328 , 0.9411)0.9411)0.9411 ). All sample size recommendations in this section pertain to the sample size per group (i.e., $n_{A}=n_{B}$). It took roughly 3 minutes to obtain this recommendation on a parallel computing server with 72 cores. While each posterior takes less than one second to approximate, we must approximate many posteriors to reliably characterize the uncertainty regarding which hypotheses are false. Our method could be expedited by considering multivariate analytical approximations to the posterior of $\boldsymbol{\theta}$. It would have taken over 20 minutes using the same computing resources to explore the sample size space using binary search instead of Algorithm 2; this discrepancy in runtime scales logarithmically with the recommended sample size.
To confirm the strong performance of our method, we estimated the joint sampling distribution of posterior probabilities for $n_{A}=13328$ under the model $\Psi$ with $m=9.9\times 10^{4}$ simulation repetitions. Using a decision threshold of $\gamma=0.9411$, the FDR and average power in (1) were estimated as 0.0497 and 0.7992. This good performance occurs under the assumption that $\boldsymbol{\eta}\sim\Psi$. If $\Psi$ is misspecified such that the effect sizes are generally overstated, the true FDR and average power are likely to exceed $q$ and not exceed $1-\beta$, respectively. In contrast, a design informed by a model $\Psi$ that generally understates the effect sizes is likely to have an FDR and average power that outperform their respective criteria. We therefore advise that practitioners define $\Psi$ with respect to the minimum effect sizes they would like to detect.
@@
To demonstrate the flexibility of our method with different models and the value of considering the dependence structure at the design stage, we implemented Algorithm 2 with a modified version of $\Psi$ where the 5 five binary outcomes are generated independently according to the marginal probabilities from a given submodel $\Psi_{\boldsymbol{\Lambda}}$. The individual metrics in $\boldsymbol{\theta}$ were also analyzed using posteriors based on independent binomial models for each outcome and diffuse BETA(1,1) priors for each marginal probability. When ignoring the dependence structure between the 5 metrics, Algorithm 2 returned a recommended $(n_{A},\gamma)$ combination of $(14427,0.9500)$ under the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. Thus, when accounting for the positive dependence between these binary outcomes, we can use a less strict decision threshold, resulting in a materially smaller sample size recommendation. If the outcomes exhibit negative dependence, we generally need a larger decision threshold – and sample size – to satisfy criteria for both the FDR and average power.
@@
To illustrate how the number of false hypotheses impacts the sample size recommendation, we implement Algorithm 2 where $\Psi$ is defined using only some of the 30 multinomial submodels. We again enforce the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. We first consider the $(n_{A},\gamma)$ recommendation where $\Psi$ is an equal mixture of the 5 submodels where only 1 hypothesis is false (i.e., where the cardinality $\lVert\boldsymbol{\Lambda}\rVert$ of the set $\boldsymbol{\Lambda}$ is 1). We separately obtain $(n_{A},\gamma)$ recommendations for the 10 submodels where $\lVert\boldsymbol{\Lambda}\rVert=2$, the 10 submodels where $\lVert\boldsymbol{\Lambda}\rVert=3$, and the 5 submodels where $\lVert\boldsymbol{\Lambda}\rVert=4$. Each recommendation was obtained via Algorithm 2 with $m=3\times 10^{4}$, $q=0.05$, and $\beta=0.2$. In Table 1, these results are compared to the $(n_{A},\gamma)$ recommendation where $\Psi$ was defined by combining all 30 multinomial submodels in an equal mixture.
$\Psi$ $n_{A}$ $\boldsymbol{\gamma}$
$\lVert\boldsymbol{\Lambda}\rVert=1$ 4627 $0.7772\times\mathbf{1}_{5}$
$\lVert\boldsymbol{\Lambda}\rVert=2$ 9985 $0.9053\times\mathbf{1}_{5}$
$\lVert\boldsymbol{\Lambda}\rVert=3$ 15035 $0.9516\times\mathbf{1}_{5}$
$\lVert\boldsymbol{\Lambda}\rVert=4$ 19668 $0.9737\times\mathbf{1}_{5}$
Combined 13328 $0.9411\times\mathbf{1}_{5}$
Combined 11106 $(0.9599,0.9599,0.9565,0.9099,0.9099)$
@@
Table 1: Recommended sample sizes $n_{A}$ and decision thresholds $\gamma$ under probability models $\Psi$ that do (Combined) and do not ($\lVert\boldsymbol{\Lambda}\rVert=\{1,2,3,4\}$) account for uncertainty in the number of false hypotheses. Two optimization schemes are considered for the Combined $\Psi$.
@@
Table 1 also demonstrates the value in considering more sophisticated optimization schemes to select $\boldsymbol{\gamma}$. For illustration, we also selected $\boldsymbol{\gamma}$ as the decision thresholds that ensure $\widehat{\text{FDR}}\leq q$ while maximizing average power subject to the constraint $\lvert\gamma_{k}-\gamma_{h}\rvert\leq 0.05$ for thresholds $k\neq h$. We detail our scheme to obtain the optimal $\boldsymbol{\gamma}_{n}$ for a given sample size $n$ in Appendix C.2. By optimizing the decision thresholds using more flexible constraints, we reduced the recommended sample size to $n_{A}=11106$. It took about 5 minutes on a parallel computing server to obtain this optimal design; 2 of these minutes were spent using a single core to implement our optimization scheme to find $\boldsymbol{\gamma}_{n}$ at each $n$ explored. While many optimization schemes not considered in this example could also be used to select $\boldsymbol{\gamma}$, the computational burden of the optimization scheme should be considered prior to implementing Algorithm 2.
@@
Lastly, we compare our sample size recommendations to those prompted by a Bonferroni correction used to control the FWER of A/B tests in frequentist settings. To bound the FWER for 5 hypotheses by 0.05, the Bonferroni correction recommends using a significance level of 0.01 for each test. This significance level is similar to a decision threshold of $\gamma=0.99$ in our context. A sample size of $n_{A}=27956$ is recommended to ensure the average power across these 5 hypothesis tests is at least 80% while bounding the FWER by 0.05. We emphasize that control of the FDR and FWER are not interchangeable. Yet by comparing this sample size recommendation to those in Table 1, we illustrate that – even with as few as $K=5$ metrics – practitioners can reduce the required sample size for their A/B tests if willing to use designs that bound the FDR instead of the FWER. The benefit of our approach will be even more pronounced in cases where $K$ is very large. While it is rare for $K$ to be very large in online controlled experiments, our method could also be applied in contexts that typically involve a much larger number of comparisons.